Compute the annual cost of each of the following expenses: $16 a week on soft drinks; $170 per month on renter's insurance. Complete the sentence: On an annual basis, the first set of expenses is _______ % of the second set of expenses.
Answer #1

He spends $16 a week on soft drinks $170 per month on renter's insurance , so the total annual expense is ,

[ there are 52 weeks and 12 months in a year ]

T= (16*52)+(170*12)= \$2872

the percentage of first set of expense or soft drink expense over the second set of expense or renter's insurance is

\frac{16*52}{170*12}*100\approx 40.784\%

So on an annual basis the first set of expenses is 40.784% of the second set of expenses